OpenStudy (yamyam70):

can someone continue the process, log(base 7) (x -5) + log (base 7 ) x + 1 = 1 log base 7 ( x-5)( x+1) = 1 7 = x^2 -4x -5 what now ?

OpenStudy (anonymous):

x^2-4x-5-7=0 x^2-4x-12=0 factorise it now

OpenStudy (yamyam70):

we have the same thought, but take a look at this.. x^2 -4x = 12 (x+2) = 12 , (x-6) = 12

OpenStudy (anonymous):

x^2-(6-2)x-12=0 can u do it now?

OpenStudy (yamyam70):

can you explain further? especially the (6-2) x

OpenStudy (anonymous):

x^2-6sx+2x-12=0 x(x-6)+2(x-6)=0 (x+2)(x-6)=0 now,x=-2 or 6

OpenStudy (yamyam70):

Thank you , instead of factoring by grouping I can equat it to zero right?

OpenStudy (anonymous):

use the process jst i did its easier.have u unrdrstood the entire process?

OpenStudy (yamyam70):

Yes :)

OpenStudy (yamyam70):

Thanks :)

OpenStudy (anonymous):

welcome